(1-p)-3 for |p| <1

Grade:

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

|p| < 1
Taylor series expansion at p = 0
(1-p)^{-3} = 1+3p+6p^2+10p^3+15p^4+21p^5+O(p^6))
Series expansion at p = 1
(1-p)^{-3} = \frac{1}{(1-p)^3}
(1-p)^{-3} = \frac{-1}{(p-1)^3}
(1-p)^{-3} = \frac{-1}{p^3-3p^3+3p-1}